# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ IV.1 Two dimensions
In our case $G=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathcal{C}[\mathbb{Z}_{2}\times\mathbb{Z}_{2}]=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. There are three generators for the SSPT. To enumerate them, let us consider a two-dimensional lattice consisting of two square sublattices. We color the sublattices with red and blue as shown in Fig. 3(a). We denote the vertices and plaquettes in each red and blue lattice by $v^{r}$ and $p^{r}$, and by $v^{b}$ and $p^{b}$, respectively. Note that the vertex of the red sublattice is effectively the plaquette of the blue sublattice and vice versa. The three generators for the SSPT are 1) $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT with cluster entangler between adjacent red and blue sites, 2) $\mathbb{Z}_{2}$ SSPT (given in section VI) on the red lattice, and 3) $\mathbb{Z}_{2}$ SSPT on the blue lattice.
@@
and similar equations for $r\leftrightarrow b$. The action of $\text{KT}^{(2)}$ on a single $Z$ operator on the red sublattice is
@@
$\textbf{D}_{r}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{\prime}\tilde{\mathbf{D}}_{x,r}H\tilde{\mathbf{D}}_{y,r}\mathbf{\tilde{P}},$ (52)
@@
$\mathbf{\tilde{P}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2}.$ (53a)
$\begin{split}\mathbf{\tilde{P}}^{\prime}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\\ &\quad\times\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (53b)
$\tilde{\mathbf{D}}_{x,r}$ and $\tilde{\mathbf{D}}_{y,r}$ in Eq. (52) denote the operator $\tilde{\mathbf{D}}_{x}$ and $\tilde{\mathbf{D}}_{y}$ on the red sublattice. $\text{KT}^{(2)}$ acting on a single $Z$ operator on the blue sublattice is
@@
$\textbf{D}_{b}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{\prime}\tilde{\mathbf{D}}_{x,b}H\tilde{\mathbf{D}}_{y,b}\mathbf{\tilde{P}},$ (56)
@@
$\begin{split}\mathbf{\tilde{P}}^{\prime}&\equiv\prod_{l=1}^{j+1}\frac{\left(1-\eta^{x}_{l+\frac{1}{2}}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l+\frac{1}{2}}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k+\frac{1}{2}}\right)}{2}\frac{(1-\eta^{y}_{i+\frac{1}{2}})}{2}\frac{(1-\eta^{y}_{i+\frac{3}{2}})}{2}\frac{(1+(-1)^{j+1}\eta^{y}_{\frac{1}{2}})}{2},\end{split}$ (57a)
$\begin{split}\mathbf{\tilde{P}}&\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l+\frac{1}{2}})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k+\frac{1}{2}})}{2}\frac{(1-\eta^{x}_{j+\frac{1}{2}})}{2}\frac{(1-\eta^{y}_{i+\frac{1}{2}})}{2}.\end{split}$ (57b)
$\tilde{\mathbf{D}}_{x,b}$ and $\tilde{\mathbf{D}}_{y,b}$ in Eq. (56) denote the operator $\tilde{\mathbf{D}}_{x}$ and $\tilde{\mathbf{D}}_{y}$ on the blue sublattice. The subsystem symmetry generators on the blue sublattice are $\eta_{j+\frac{1}{2}}^{x}=\prod_{i=0}^{L_{x}-1}X_{i+\frac{1}{2},j+\frac{1}{2}}$ and $\eta_{i+\frac{1}{2}}^{y}=\prod_{j=0}^{L_{y}-1}X_{i+\frac{1}{2},j+\frac{1}{2}}$. The structure of the membrane operator in Eq. (50) and Eq. (54) (see Fig. 4b & c) for the derived $\text{KT}^{(2)}$ transformation resembles the structure given in Ref. [28], except for the rotation of the lattice by $45^{\circ}$. Their mapping of a Pauli $Z$ operator involves a product of Pauli operators supported on a quadrant (light cone) relative to the original Pauli $Z$ operator, where the original ones on the red and blue lattices get mapped to the opposite quadrants. Similar to theirs, we find a light cone structure in our transformation: We also find that the composition of two $\text{KT}^{(2)}$ transformations gives
$(\text{KT}^{(2)})^{2}=\mathbf{P}_{r}\mathbf{P}_{b},$ (58)
$\mathbf{P}_{r}$ and $\mathbf{P}_{b}$ are the respective projections onto the Hilbert spaces of the red and blue sublattices, which are separately subsystem symmetric.
@@ IV.2 Higher dimensions
Here, we will generalize all the discussions we had for two dimensions to higher dimensions. We consider $d$ dimensional lattice consisting of two hypercubic sublattices. These sublattices are dual to each other. Again, we colour them by red and blue. Let us denote the vertices and hypercubes of the red and blue sublattices by $v^{r}$,$c^{r}$ and $v^{b}$,$c^{b}$ respectively. A vertex of red sublattice is effectively a hypercube of the blue sublattice. Let us consider the $d$-dimensional entangler on the red and blue hypercubic sublattices for the cluster state described by the Hamiltonian
@@
By $r\leftrightarrow b$, we get another similar set of equations. The action of $\text{KT}^{(d)}$ on a single $Z$ operator can be worked out similarly in the two-dimensional case, and we find that the light cone structure is in opposite directions on the two sublattices [8]. The composition of two $\text{KT}^{(d)}$ transformations is
$(\text{KT}^{(d)})^{2}=\mathbf{P}_{r}^{(d)}\mathbf{P}_{b}^{(d)}.$ (62)
$\mathbf{P}_{r}^{(d)}$ and $\mathbf{P}_{b}^{(d)}$ are projections onto subsystem symmetric subspace of the Hilbert spaces of red and blue sublattices.
@@ V.1 Two dimensions
$\mathbf{D}_{SSPT}^{(2)}\equiv\text{KT}^{(2)\dagger}\>\mathbf{D}^{(2)}\>\text{KT}^{(2)}.$ (63)
@@
$\left(\mathbf{D}_{SSPT}^{(2)}\right)^{2}=\mathbf{T}_{(1,1)}\mathbf{P}.$ (65)
@@ V.2 Higher dimensions
$\mathbf{D}_{SSPT}^{(d)}\equiv\text{KT}^{(d)\dagger}\>\mathbf{D}^{(d)}\>\text{KT}^{(d)}.$ (66)
@@
where $d$ denotes the direction $(1,1,...,1)$ and is the diagonal vector. The vertices $v^{r}$ and $v^{b}$ in Eq. (67a) are related by a half diagonal translation $(\frac{1}{2},\frac{1}{2},...,\frac{1}{2})$, i.e., $v^{b}=v^{r}+(\frac{1}{2},\frac{1}{2},...,\frac{1}{2})$. Hence, $\mathbf{D}_{SSPT}^{(d)}$ is a symmetry of the d-D cluster state Hamiltonian Eq. (59) with $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ subsystem symmetry. Again, it squares to diagonal translation times a projection,
$(\mathbf{D}_{SSPT}^{(d)})^{2}=\mathbf{T}_{(1,...,1)}\mathbf{P}.$ (68)
{tikzpicture}

### v2

@@ IV.1 Two dimensions
As a remark by [9], this is the classification of strong SSPT phases with symmetry group $G$. There is also a notion of weak SSPT phases which is composed of decoupled 1D SPTs. We will not consider weak SSPT phases in our analysis. In our case $G=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathcal{C}[\mathbb{Z}_{2}\times\mathbb{Z}_{2}]=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. There are three generators for the SSPT. To enumerate them, let us consider a two-dimensional lattice consisting of two square sublattices. We color the sublattices with red and blue as shown in Fig. 4(a). We denote the vertices and plaquettes in each red and blue lattice by $v^{r}$ and $p^{r}$, and by $v^{b}$ and $p^{b}$, respectively. Note that the vertex of the red sublattice is effectively the plaquette of the blue sublattice and vice versa. The three generators for the SSPT are 1) $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT with cluster entangler between adjacent red and blue sites, 2) $\mathbb{Z}_{2}$ SSPT (given in section VI) on the red lattice, and 3) $\mathbb{Z}_{2}$ SSPT on the blue lattice.
@@
and similar equations for $r\leftrightarrow b$. Readers may wonder definitions of the order parameter of the SSSB phase and the string order parameter, and how they are related by the KT transformation. We discuss this in Appendix G. The action of $\text{KT}^{(2)}$ on a single $Z$ operator on the red sublattice is
@@
$\textbf{D}_{r}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}\tilde{\mathbf{D}}_{x,r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}_{r}}\tilde{\mathbf{D}}_{y,r}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}},$ (54)
@@
$\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2}.$ (55a)
$\begin{split}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\\ &\quad\times\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (55b)
$\tilde{\mathbf{D}}_{x,r}$ and $\tilde{\mathbf{D}}_{y,r}$ in Eq. (54) denote the operator $\tilde{\mathbf{D}}_{x}$ and $\tilde{\mathbf{D}}_{y}$ on the red sublattice and $\mathbf{H}^{\otimes(2)}_{r}$ denote the tensor product of Hadamard operator on all sites of the red sublattice. $\text{KT}^{(2)}$ acting on a single $Z$ operator on the blue sublattice is
@@
$\textbf{D}_{b}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}\tilde{\mathbf{D}}_{x,b}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}_{b}}\tilde{\mathbf{D}}_{y,b}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}},$ (58)
@@
$\begin{split}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}&\equiv\prod_{l=1}^{j+1{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1}}\frac{\left(1-\eta^{x}_{l+\frac{1}{2}}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l+\frac{1}{2}}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k+\frac{1}{2}}\right)}{2}\frac{(1-\eta^{y}_{i+\frac{1}{2}})}{2}\frac{(1-\eta^{y}_{i+\frac{3}{2}})}{2}\frac{(1+(-1)^{j+1}\eta^{y}_{\frac{1}{2}})}{2},\end{split}$ (59a)
$\begin{split}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}&\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l+\frac{1}{2}})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k+\frac{1}{2}})}{2}\frac{(1-\eta^{x}_{j+\frac{1}{2}})}{2}\frac{(1-\eta^{y}_{i+\frac{1}{2}})}{2}.\end{split}$ (59b)
$\tilde{\mathbf{D}}_{x,b}$ and $\tilde{\mathbf{D}}_{y,b}$ in Eq. (58) denote the operator $\tilde{\mathbf{D}}_{x}$ and $\tilde{\mathbf{D}}_{y}$ on the blue sublattice and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}_{b}}$ denote the tensor product of Hadamard operator on all sites of the blue sublattice. The subsystem symmetry generators on the blue sublattice are $\eta_{j+\frac{1}{2}}^{x}=\prod_{i=0}^{L_{x}-1}X_{i+\frac{1}{2},j+\frac{1}{2}}$ and $\eta_{i+\frac{1}{2}}^{y}=\prod_{j=0}^{L_{y}-1}X_{i+\frac{1}{2},j+\frac{1}{2}}$. The structure of the membrane operator in Eq. (52) and Eq. (56) (see Fig. 4b & c) for the derived $\text{KT}^{(2)}$ transformation resembles the structure given in Ref. [32], except for the rotation of the lattice by $45^{\circ}$. Their mapping of a Pauli $Z$ operator involves a product of Pauli operators supported on a quadrant (light cone) relative to the original Pauli $Z$ operator, where the original ones on the red and blue lattices get mapped to the opposite quadrants. Similar to theirs, we find a light cone structure in our transformation. We also find that the composition of two $\text{KT}^{(2)}$ transformations gives
$(\text{KT}^{(2)})^{2}\propto\,\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{r}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{b},$ (60)
$\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{r}$ and $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{b}$ are the respective projections onto the subspaces of the Hilbert spaces of the red and blue sublattices, which are separately subsystem symmetric.
@@ IV.2 Higher dimensions
Here, we will generalize all the discussions we had for two dimensions to higher dimensions. We consider $d$ dimensional lattice consisting of two hypercubic sublattices. These sublattices are dual to each other. Again, we color them red and blue. Let us denote the vertices and hypercubes of the red and blue sublattices by $v^{r}$,$c^{r}$ and $v^{b}$,$c^{b}$ respectively. A vertex of the red sublattice is effectively a hypercube of the blue sublattice. Let us consider the $d$-dimensional entangler on the red and blue hypercubic sublattices for the cluster state described by the Hamiltonian
@@
By $r\leftrightarrow b$, we get another similar set of equations. The action of $\text{KT}^{(d)}$ on a single $Z$ operator can be worked out similarly in the two-dimensional case, and we find that the light cone structure is in opposite directions on the two sublattices [8]. The composition of two $\text{KT}^{(d)}$ transformations is proportional to a projection
$(\text{KT}^{(d)})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{P}_{r}^{(d)}\mathbf{P}_{b}^{(d)}.$ (64)
$\mathbf{P}_{r}^{(d)}$ and $\mathbf{P}_{b}^{(d)}$ are projections onto subsystem symmetric subspace of the Hilbert spaces of red and blue sublattices. In appendix F, we describe how to implement the KT transformation in a short-depth operation.
@@ V.1 Two dimensions
$\mathbf{D}_{SSPT}^{(2)}\equiv\text{KT}^{(2)\dagger}\>\mathbf{D}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}^{(2)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{D}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}^{(2)}}\>\text{KT}^{(2)}.$ (65)
@@
$\left(\mathbf{D}_{SSPT}^{(2)}\right)^{2}\propto\,\mathbf{T}_{(1,1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},r}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{(1,1),b}}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}}.$ (67)
The translations $\mathbf{T}_{(1,1),r}$ and $\mathbf{T}_{(1,1),b}$ act on both the red and blue sublattice by a diagonal unit shit on the respective sublattices. We note that the transformations Eq. (66) are equivalently achieved by the controlled-Z operator $\mathcal{U}_{CZ}^{(2)}=\prod_{v^{r}}\prod_{v^{b}=p^{r}\in\partial^{*}v_{r}}CZ_{v^{r},p^{r}}$ with $\partial^{*}v_{r}$ a set of plaquettes that contain $v_{r}$. $\mathcal{U}_{CZ}^{(2)}$ is a unitary and squares to 1.
@@ V.2 Higher dimensions
$\mathbf{D}_{SSPT}^{(d)}\equiv\text{KT}^{(d)\dagger}\>\mathbf{D}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}^{(d)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{D}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}^{(d)}}\>\text{KT}^{(d)}.$ (68)
@@
where $d$ denotes the direction $(1,1,...,1)$ and is the diagonal vector. The vertices $v^{r}$ and $v^{b}$ appearing as subscripts of Pauli $X$ operators in Eq. (69) are related by a half diagonal translation $(\frac{1}{2},\frac{1}{2},...,\frac{1}{2})$, i.e., $v^{b}=v^{r}+(\frac{1}{2},\frac{1}{2},...,\frac{1}{2})$. Hence, $\mathbf{D}_{SSPT}^{(d)}$ is a symmetry of the d-D cluster state Hamiltonian Eq. (61) with $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ subsystem symmetry. Again, it squares to diagonal translation times a projection up to a normalization factor
$(\mathbf{D}_{SSPT}^{(d)})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{T}_{(1,...,1),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{(1,...,1),b}}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b}}}.$ (70)
The translations $\mathbf{T}_{(1,...,1),r}$ and $\mathbf{T}_{(1,...,1),b}$ act on both the red and blue sublattice by a diagonal unit shit on the respective sublattices. We note that the transformations Eq. (69) are equivalently achieved by the controlled-Z operator $\mathcal{U}_{CZ}^{(d)}=\prod_{v^{r}}\prod_{v^{b}=c^{r}\in\partial^{*}v_{r}}CZ_{v^{r},c^{r}}$ with $\partial^{*}v_{r}$ a set of hypercubes that contain $v_{r}$. $\mathcal{U}_{CZ}^{(d)}$ is a unitary and square to 1.